Radius of gyration of disc rotating about an axis perpendicular to its passing through its centre is_____?

Grade:12

1 Answers

Vikas TU
3 years ago
I = MK^2
 
↪so we get , the Radius of Gyration as ,
 
〽K = underoot (I / M)
 
↪So now here for getting the Radius of Gyration of an object across the diameter , we apply here the perpendicular Axis theorem
 
↪Iz = Ix + Iy
 
↪since , here , we get symmetry along all axis , we get as ,
 
↪( Ix = Iy)
 
↪Iz = 2 Ix
 
↪Ix = Iz/2
 
↪so now here, we apply , we get as
 
⭐K' = root 2 cm